OpenStudy (anonymous):

Using substitution, could someone please help me with this system? –y + 3x = 6 y = –6x + 12 So far I've gotten: Substituting the second equation into the first you get -(-6x+12) + 3x = 6 Which can be solved for x after finding x you can use the second equation above directly to find y

OpenStudy (anonymous):

Sounds like you know what you are doing .... What's the problem you are having?

OpenStudy (anonymous):

Well, I had help with that part, but I can't figure out what to do next. I tried using my textbook, but it was a bit confusing, because they showed you what to do, but didn't explain it.

OpenStudy (anonymous):

Ok, first let's distribute that -(-6x+12) Can you do that?

OpenStudy (anonymous):

-6x + -72?

OpenStudy (anonymous):

Think of it this way: (-)(-6x) & (-)(12)

OpenStudy (anonymous):

Because negative times negative equals a positive 6x + -12?

OpenStudy (anonymous):

yes, very good so we have 6x-12+3x=6 What will you do next?

OpenStudy (anonymous):

Add 6x and 3x?

OpenStudy (anonymous):

correct 9x-12=6 now what?

OpenStudy (anonymous):

Substitute for y from the other equation?

OpenStudy (anonymous):

Let's try to solve for x.

OpenStudy (anonymous):

Sorry, I was looking at my textbook, and thought that we were onto that step.

OpenStudy (anonymous):

:-) That's ok.

OpenStudy (anonymous):

Okay, so how would I go about solving for x? Also, thank you for the medal.

OpenStudy (anonymous):

We have x on one side, can we get non x-values on the other side?

OpenStudy (anonymous):

You would add 12 + 6? Then get 18, divided by 9x would equal 2.

OpenStudy (anonymous):

Excellent: So x=2

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

What are you going to do with that x=2?

OpenStudy (anonymous):

You still there?

OpenStudy (anonymous):

Substitute 2 for x in either equation and solve for y? Sorry, my computer is slow sometimes.

OpenStudy (anonymous):

Heh I Understand. You are correct, good job!

OpenStudy (anonymous):

Thank you! So, -y + 3(2) + 6.

OpenStudy (anonymous):

-y+3(2)=6

OpenStudy (anonymous):

So, -y + 6 = 6.

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

What would I do next?

OpenStudy (anonymous):

same that you did when you had 9x-12=6 and solved for x. now we are going to it for y

OpenStudy (anonymous):

6 - 6 = 0 So y, would be 0?

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

Then I would substitute 0for y, and get 0 + 6 = 6.

OpenStudy (anonymous):

Correct: So your Answer is: X=2, y=0

OpenStudy (anonymous):

As long as you get a correct statement when you plug them both back in, like you showed 6=6 and not get something like 7=6, then we would have a problem.

OpenStudy (anonymous):

Okay, thank you some much for your help. If I could give you more than one medal I would.

OpenStudy (anonymous):

Thank you, It is nice to meet someone who wants to work out the problems and not just demand an answer.

OpenStudy (anonymous):

I get what you mean, you are a very good teacher by the way.